OpenStudy (anonymous):

x-y+z=10 4x+y-2z=15 -3x+5y-z=-18 This is a Linear Equation With Three Variables and I must find x y and z

OpenStudy (anonymous):

Solve for one variable and use it in the second eq. x = 10 + y - z use it in second 4(10 + y - z) + y - 2z = 15 simplify 40 + 4y - 4z + y - 2z = 15 5y - 6z = -35 solve this for y -7 + 6/5 z = y Now you have values for both x and y. You can use them to write a complete function in terms of z, and then solve for the z value. Repeat this procedure and you can get values for each variable

OpenStudy (anonymous):

Thank you so much.